3. Find the time function corresponding to the following Laplace Transform. You must solve by hand using the inverse LaplaceI just need answer for part b.
Answer #1

325 s 3+2 F (S) 6y banta rahin AS+B 32s st ) 32544 3+254= (aS+6) (s- 16)+ cs-4s-16 C : (As3ASB-i<] cs3-4cs24cs 6 C DS3 +4p2 +wnty eu aho and , -16A4C40 = 2 dolvi abo Rahoy A 0.1 f32 d-53 125 6 36 8 0 CCD) = x [-) Also B 2a 59//6a f17/s2 4(FIS)) To 20